### v1

Abstract
We explore the application of stabilizer formalism in deriving the two Bloch Sphere representation of pure two-qubit states, including maximally entangled ones. Specifically, we investigate the stabilizer groups of two-qubit stabilizer states and establish a link with both the density matrix and graphical representation of a given stabilizer state. The sum of the stabilizer group’s matrices constitutes the state’s density matrix. Simultaneously, matrices of a stabilizer group of a given state may be associated with the coordinate axes of the Bloch Spheres representing that state. This analysis reveals a natural link between the density matrix and graphical representations, offering insights into geometry and symmetries of entanglement and two-qubit states in general.
I 1. Introduction
Bloch Sphere representation [1, 2, 3, 4] is a powerful visualization tool that gets used across various areas of Quantum Mechanics. Extension of this representation to systems of two or more qubits is challenging due to entanglement. Recently there appeared a way to represent any pure two-qubit state using two Bloch Spheres [5]. The present work builds up on that development. We take a step back and focus on connection between a stabilizer group of a given two-qubit stabilizer state and the graphical representation of that state using two Bloch Spheres.
The work is organized as follows. In Chapter 2 we briefly introduce stabilizer states and stabilizer groups. In Chapter 3 we speak about connection between the stabilizer group of a given state and its density matrix. In Chapter 4 we briefly introduce the graphical representation developed in [5] with focus on stabilizer state representation. In Chapter 5 we speak about connection between the stabilizer group of a given state and its graphical representation. As a result we obtain the connection between a density matrix of a given state and its representation on two Bloch Spheres. In Chapter 6 we discuss why and how the correspondence works. And chapter 7 is Conclusion.
II 2. Two-Qubit Stabilizer States
We use this chapter to introduce stabilizer states and stabilizer groups so that the paper is self-contained. We do not attempt an exhaustive exposition of the topic, instead we do the introduction that makes it easy to understand the rest of the paper. For a rigorous account we refer the reader to [6] Chapter 10 or [7].
Essentially, two-qubit stabilizer states are those states that are reachable from the state $\ket{\uparrow\uparrow}$ by Clifford gates (Hadamard; Phase; CNOT). Put differently, let us take the set of $4\times 4$ two-qubit Pauli matrices that have a form $I\otimes P$ or $P\otimes I$ or $P\otimes P$ where $I$ is $2\times 2$ Identity and P is some $2\times 2$ Pauli matrix $X,Y,Z$. Now let us navigate the space of pure two-qubit states by repeatedly applying these matrices to the state $\ket{\uparrow\uparrow}$ and restricting the angle of rotation to $90$ degrees. The states reachable in this way are the two-qubit stabilizer states.
There are 60 two-qubit stabilizer states: 36 separable and 24 maximally entangled ones [8, 9]. Stabilizer group of a given state is formed by matrices that, when applied to the given state, leave it unchanged. Essentially the members of the stabilizer group, besides Identity, are eigenmatrices of a given state times the eigenvalue. Below are correspondences between some of the stabilizer states and their stabilizer groups. We have chosen four states, two separable and two entangled: $\ket{\uparrow\uparrow}$; |$\ket{\uparrow\leftarrow}=\frac{1}{\sqrt{2}}(\ket{\uparrow\uparrow}-\ket{\uparrow\downarrow})$; |$\ket{\Psi^{-}}=\frac{1}{\sqrt{2}}(\ket{\uparrow\downarrow}-\ket{\downarrow\uparrow})$ and |$\ket{\Phi^{+i}}=\frac{1}{\sqrt{2}}(\ket{\uparrow\uparrow}+i\ket{\downarrow\downarrow})$
$\ket{\uparrow\uparrow}\leftrightarrow$$\{I\otimes I;Z\otimes I;I\otimes Z;Z\otimes Z\}$ (1)
$\ket{\uparrow\leftarrow}\leftrightarrow$$\{I\otimes I;Z\otimes I;-I\otimes X;-Z\otimes X\}$
|$\ket{\Psi^{-}}\leftrightarrow$$\{I\otimes I;-Z\otimes Z;-X\otimes X;-Y\otimes Y\}$
|$\ket{\Phi^{+i}}\leftrightarrow$$\{I\otimes I;Z\otimes Z;X\otimes Y;Y\otimes X\}$
Stabilizer group of a pure two-qubit state consists of 4 matrices, one of which is always $4\times 4$ Identity. Separable state has always two matrices involving tensor product of a Pauli matrix with Identity and entangled state has three matrices where the tensor product happens between two Pauli matrices.
III 3. Connecting stabilizer group and density matrix
The state’s density matrix may be related to the stabilizer group of a state in a straightforward way. One just needs to sum the matrices that make up the stabilizer group of a state. More formally, given n-qubit state |$\ket{\Psi}$ and its stabilizer group $\mathcal{S}$ consisting of $2^{n}$ matrices $M\in\mathcal{S}$ s.t. $M\ket{\Psi}=\ket{\Psi}$ for all $M$ in $\mathcal{S}$ the following is true [10]:
$\rho_{\mathcal{S}}=\ket{\Psi}\bra{\Psi}=\frac{1}{2^{n}}\sum\limits_{M\in\mathcal{S}}M$ (2)
Having linked the density matrix of a state to its stabilizer group we may turn to linking its stabilizer group to graphical representation.
IV 4. Representing stabilizer states within two Bloch Sphere formalism
In order to make the aforementioned connection we need to briefly describe the method of representation of pure two-qubit states on two Bloch spheres introduced in [5]. In this work we are not concerned with partially entangled states because two-qubit stabilizer states are either separable or maximally entangled. Hence, although in the original paper the partially entangled state representation is also described, we will not touch upon it here. For detailed analysis of graphical representations described further we refer the reader to the aforementioned paper.
Representation of separable two-qubit stabilizer states is simple. The statevector arrow of each Bloch sphere is facing in one of the six directions: ±$\pm x;\pm y;\pm z$. There are $6\times 6=36$ configurations in total. See Figure 1 for an example.
$x_{1}$$y_{1}$$z_{1}$
$x_{2}$$y_{2}$$z_{2}$
Figure 1: Two Bloch Sphere representation of the separable state $\ket{\uparrow\uparrow}$
Representation of the entangled two-qubit stabilizer states goes as follows. There are no statevector arrows - instead there is just a dot in the middle of each BS, i.e. the statevector arrows are facing in the $(0,0,0)$ direction. The coordinate axes of the Bloch Spheres have different handedness and the relative direction of the coordinate axes encodes the state. Therefore, having fixed the coordinates of the first BS, the $z_{2}$ axis may face any of the 6 directions; for each of those $x_{2}$ axis may face 4 possible directions because it has to also be orthogonal to $z_{2}$; the combination of $z_{2}$ and $x_{2}$ defines the direction of $y_{2}$ unambiguously. Therefore there are $6\times 4=24$ possible configurations. See Figure 2 for an example.
$x_{1}$$y_{1}$$z_{1}$
$x_{2}$$y_{2}$$z_{2}$
Figure 2: Two Bloch Sphere representation of the entangled state |$\ket{\Psi^{-}}$
Representing entanglement through the relative alignment of coordiante axes of the two Bloch Spheres makes it obvious why the state of single qubit of the entangled pair is maximally mixed and why we need to have access to the whole system to extract any information about the two-qubit state.
Furthermore, we can define rotations that don’t involve change of direction of the statevector. Namely, we can use passive rotations - or rotations of coordinate axes - to depict unitary evolution of the system. Although the local rotations refer to coordinate rotations on a single BS, if we start from a maximally entangled state and restrict ourselves to only local rotations of a single qubit - we can still reach any relative alignment of the coordinate axes and therefore any maximally entangled state. The representation in Figure 2 clarifies this quantum mechanical fact.
Direct consequence of adopting passive rotations is the immovability of the statevector arrows. In fact, we can always choose the statevector arrows to point in the same direction as is done in Figure 1. Such convention will be useful for graphical interpretation of the stabilizer group. Having introduced the essentials of two Bloch Sphere geometry of two-qubit states we can proceed to connect it to the stabilizer group structure.
V 5. Connecting stabilizer group to graphical representation
Let us examine the stabilizer groups of the states $\ket{\uparrow\uparrow}$ and |$\ket{\Psi^{-}}$ in equation 1 and Bloch Sphere representations of those states in Figures 1 and 2.
The representation of the state |$\ket{\Psi^{-}}$ has the coordinate axes of both BS always facing in the opposite direction. The respective stabilizer group is $\{I\otimes I;-Z\otimes Z;-X\otimes X;-Y\otimes Y\}$. We could interpret members of the group that have a form $-Z\otimes Z$ as an indication of the relative direction of the respective coordinate axes. In other words we can say that $z_{1}$ axis is facing in the same direction as $-z_{2}$ axis. Note that nothing would change if we were to assign the -- sign to $z_{1}$ axis and say that $-z_{1}$ is facing in the same direction as $z_{2}$. Interpreting other members of the group in the same way we arrive at the representation from Figure 2.
The stabilizer group of the state $\ket{\uparrow\uparrow}$ is $\{I\otimes I;Z\otimes I;I\otimes Z;Z\otimes Z\}$. Along with the member $Z\otimes Z$ which we know how to interpret, it has members of the form $Z\otimes I$. We propose to interpret structure like this as indicating which axis is facing along the statevector arrow. In this case $z_{1}$ axis is facing along the statevector arrow. In the same way we arrive at the interpretation of $I\otimes Z$: $z_{2}$ axis is facing along the statevector arrow. We have chosen the statevector arrows to remain static, therefore we define interpretations in terms of coordinate axes.
Let us now use the aforementioned rules to create representations of the rest of the states mentioned in eq. 1. The stabilizer group of the state $\ket{\uparrow\leftarrow}$ is $\{I\otimes I;Z\otimes I;-I\otimes X;-Z\otimes X\}$. $Z\otimes I$ Implies that the axis $z_{1}$ will face along the statevector arrow; $-I\otimes X$ implies that the axis $-x_{2}$ will be facing along the statevector arrow; $-Z\otimes X$ implies that the axis $z_{1}$ will be facing in the same direction as axis $-x_{2}$. Figure 3 contains the representation. Indeed that is a representation of the state $\ket{\uparrow\leftarrow}$.
$x_{1}$$y_{1}$$z_{1}$
$z_{2}$$y_{2}$$x_{2}$
Figure 3: Two Bloch Sphere representation of the separable state $\ket{\uparrow\leftarrow}$
The stabilizer group for the state |$\ket{\Phi^{+i}}$ is $\{I\otimes I;Z\otimes Z;X\otimes Y;Y\otimes X\}$. This state is maximally entangled, therefore there will be no statevector arrows. Interpretation of the three members of the group gives us the following pairs of co-directional axes: $\{z_{1}\And z_{2}$; $x_{1}\And y_{2}$; $y_{1}\And x_{2}\}$ resulting in the representation shown in Figure 4. Note that the second BS has left-handed coordinates.
$x_{1}$$y_{1}$$z_{1}$
$y_{1}$$x_{1}$$z_{1}$
Figure 4: Two Bloch Sphere representation of the entangled state |$\ket{\Phi^{+i}}$
Linking stabilizer group to graphical representation on one hand and to density matrix on the other - we effectively link the latter two. In the next chapter we will speak about rotations to explain how the movement between the states happens, show some useful applications of the two BS formalism and reflect on why this construction is working.
VI 6. Why and how this works.
When we speak about movement between states we speak of rotation. Usually rotation is defined to happen around some axis, but it can also be defined to happen in some plane. It has been shown [12] that rotation generated by one of 15 $4\times 4$ two-qubit Pauli matrices can be associated with rotation in one of 15 planes formed by a pair of axes from the set $\{z_{1};x_{1};y_{1};z_{2};x_{2};y_{2}\}$. For the purpose of simplicity we will only speak about planes formed by axes of the same Bloch Sphere, i.e. local rotations. For an extended discussion of arbitrary rotations in arbitrary plane we refer the reader to [5].
To move between the states we apply to an initial state a Unitary of the form $e^{i\frac{\theta}{2}(P)}$ where $P$ is one of 15 two-qubit Pauli matrices and $\theta$ is the angle of rotation. This is equivalent to a movement of coordinate axes. For example to move from state $\ket{\uparrow\uparrow}$ to the state $\ket{\uparrow\leftarrow}$ we apply $e^{i\frac{\pi}{4}(I\otimes\sigma_{y})}$ to the former. This is equivalent to rotation of the coordinate plane formed by the axes $z_{2}$ and $x_{2}$ by $\pi/2$ (compare figures 1 and 3).
In [5] it has been shown how having graphical representation of states and unitaries helps to understand quantum gates, for example CNOT. Here we would like to briefly discuss another advantage that requires understanding only of local rotations. Let us look at the Figure 2. Let’s say we are interested in exploring the symmetry of the state |$\ket{\Psi^{-}}$. We want to find some other maximally entangled state that is related to |$\ket{\Psi^{-}}$ in an interesting way. One such state might be the following state: the coordinate axes of the second BS representation are cyclically permuted (see Figure 5).
$x_{1}$$y_{1}$$z_{1}$
$z_{2}$$x_{2}$$y_{2}$
Figure 5: Two Bloch Sphere representation of the entangled state obtained by cyclic permutation of second BS coordinate axes of the state |$\ket{\Psi^{-}}$$\{x_{2};y_{2};z_{2}\}$ to $\{z_{2};x_{2};y_{2}\}$
From the visual representation we infer that this state is related to |$\ket{\Psi^{-}}$ by $\theta=2\pi/3$ rotation of the second BS in the plane perpendicular to the axis $z_{2}+x_{2}+y_{2}$. The stabilizer group of the state can also be inferred from the representation: $\{I\otimes I;-Z\otimes Y;-Y\otimes X;-X\otimes Z\}$. Guided by equation 2 we can sum the group members to obtain density matrix (eq. 3).
$\rho=\frac{1}{4}\begin{pmatrix}1&i&-1&i\\ -i&1&i&1\\ -1&-i&1&-i\\ -i&1&i&1\end{pmatrix}$ (3)
From the density matrix we can obtain the ket form of the state: $\frac{1}{2}\ket{\uparrow\uparrow-i\uparrow\downarrow-\downarrow\uparrow-i\downarrow\downarrow}$. We can do another cycle of permutations on the second BS and obtain the graphical representation, stabilizer group, density matrix and ket of a state $\frac{1}{2}\ket{\uparrow\uparrow-\uparrow\downarrow-i\downarrow\uparrow-i\downarrow\downarrow}$. Looking only at the ket forms of the three states one would hardly see that they are related by a simple symmetry.
Next we would like to discuss an observation that might provide a hint to why relations discussed in the paper take place. Let us look at the Figure 5 and the stabilizer group of the corresponding state $\{I\otimes I;-Z\otimes Y;-Y\otimes X;-X\otimes Z\}$.
If we hold the axes $x_{1}$ and $-z_{2}$ and rotate the Bloch Spheres simultaneously around those axes we keep the relative alignment of all the axes. Remembering that it is the relative alignment of the axes that determines the state, the state hasn’t been changed by rotation. Therefore we can say that the stabilizer group indicates the pairs of axes simultaneous rotation about which keeps the state unchanged.
For the case of separable states the interpretation is similar. Let us take the state $\ket{\uparrow\leftarrow}$ with stabilizer group $\{I\otimes I;Z\otimes I;-I\otimes X;-Z\otimes X\}$ depicted in Figure 3. $Z\otimes I$ means that rotation about the $z_{1}$ axis doesn’t change the state, same for the $-x_{2}$ axis. Furthermore the simultaneous rotation about $z_{1}$ and $-x_{2}$ axis also leaves the state unchanged. Effectively, the group members of the form $I\otimes P$ link rotation around a given axis with rotation around the statevector arrow, i.e. highlighted direction in space.
The last observation we would like to make in this work is that in the representation of the separable state we know only about the axis that is facing in the direction of the statevector. For example in Figure 1 we are actually ignorant about where the axes $x_{1},y_{1},x_{2},y_{2}$ are facing and we have drawn the axes in this way to have familiar coordinate structure. The axes could be facing anywhere in the plane perpendicular to the statevector arrow as long as the axes $x$ and $y$ are mutually perpendicular. We could say that in separable mode there is a highlighted direction, direction of statevector arrows, and we know only about the axes that are facing in that direction. In fact, this ambiguity is important in movement between separable and entangled modes [5].
VII 7. Conclusion
In this work we have been developing the graphical representation of pure two-qubit states on two Bloch Spheres. We have shown a way to connect algebraic state form (ket vector, density matrix) with the graphical representation. This is done via the stabilizer group of the state. Stabilizer group translates in a straightforward way into both algebraic and graphical form. To obtain density matrix from a stabilizer group one just needs to sum the group members. To obtain the graphical representation one just needs to interpret the group members as co-aligned coordinate axes.
Although in this work the correspondence is shown only for stabilizer states, it can be extended to arbitrary pure states as any pure state may be expressed through stabilizer states. Moreover, the graphical method discussed here is also built on stabilizer state representations, which are used as building blocks to represent arbitrary pure states.
This work brings us closer to two Bloch Sphere visualization that is intuitive and easy to manipulate. Already as they are, the correspondences described in this work allow for exploration of non-trivial symmetries in two-partite Hilbert space and visual analysis of quantum circuits.
The correspondence between stabilizer group and graphical representation might be understood if one interprets Bloch Sphere as a visualization of what kinds of rotation leave the state unchanged.
On the level of quantum foundations the aforementioned correspondence reiterates and solidifies the observation that entanglement is related to inversion of particle’s internal coordinates.
Acknowledgements.